OpenStudy (drummmlers):

The hypotenuse of a right triangle is 10 inches. The shorter leg is 10 inches less than twice the longer leg. Find the lengths of the two legs. This is what I have so far (2x - 10)^2 +(x)^2 =10^2

OpenStudy (anonymous):

Ok Let us take the length of the longer leg as x So the length of the shorter leg would be 2x-10 So, according to pythagoras theorem, (2x - 10)^2 +(x)^2 =10^2 4x^2 + 100 - 40x + x^2 = 100 5x^2 - 40x = 0 So x can assume two values. 0 or 8. But, the side of a triangle can't be 0 So , length of longer leg would be 8 length of shorter leg would be 16-10 = 6 Hope you're satisfied with the answer